OpenStudy (anonymous):

Y varies directly as to x. When x = -8 y = 6. Find y when x = 10

OpenStudy (mayankdevnani):

\[\huge \bf y=k \times x\] where k is any number(resultant) so, \[\huge \bf k=\frac{y}{x}\] If x=-8 and y=6, \[\huge \bf k=\frac{6}{-8}=\frac{-3}{4}\]

OpenStudy (mayankdevnani):

If x=10 and y=?, \[\huge \bf k=\frac{y}{x}=\frac{y}{10}\] \[\huge \bf \frac{-3}{4}=\frac{y}{10}\] now solve for y.

OpenStudy (mayankdevnani):

can you solve it? @balokkk

OpenStudy (anonymous):

so the answer is -7.5? @mayankdevnani

OpenStudy (mayankdevnani):

\[\huge \bf 4y=-30\] By cross multiply ^^^ divide 4 on both the sides, \[\huge \bf y=\frac{-30}{4}=-7.5\] you are correct @balokkk

OpenStudy (mayankdevnani):

and also you can give me medal by clicking BEST RESPONSE right side to my response.

OpenStudy (mayankdevnani):

good. @balokkk